Two planes leave the same airport the same time and travel in opposite directions, one at 550 miles per hour and the other at 300 miles per hour. In how many hours will they be 2550 miles apart?
Answer by TimothyLamb(4379)

s = d / t
t = d / s
t = 2550 / (550 + 300)
Answer:
t = 3 hours
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php